OpenStudy (anonymous):

6. Mr. Crisostomo is 30 years older than his son. How old are the father and the son if Mr. Crisostomo is 10 more than thrice as old as his son?

terenzreignz (terenzreignz):

The key to this kind of question is knowing how to represent values. If we let x be the age of Mr. Crisostomo's son, how old is Mr. Crisostomo himself?

terenzreignz (terenzreignz):

HINT: Use the fact that Mr. Crisostomo is \(\color{red}{\text{30 years older}}\) than his son.

OpenStudy (anonymous):

done :D

terenzreignz (terenzreignz):

That was fast. What's your answer?

OpenStudy (anonymous):

Let x = Mr. Crisostomo’s Age Let y = Son’s Age X = 30 + Y How if: X = 3Y + 10 30 + Y = 3Y + 10 30 – 10 = 3Y – Y 20 = 2Y 20/2 = 2Y 10 = Y The Son is 10 years old X = 30 + Y X = 30 + 10 X = 40 Mr. Crisostomo’s Age = 40 any correctios would be gladly accepted :D

OpenStudy (anonymous):

**corrections

terenzreignz (terenzreignz):

Fair enough. So... WHY did you post this when you could do it yourself? XD #GalingNamanNiKuya :P

OpenStudy (anonymous):

uhh.. just to have a helped feeling :D,,, finishing five thesis is the worst thing i have to experience this day D: anyway Thanks :DD

terenzreignz (terenzreignz):

But I did virtually nothing... anyway, stick around XD

OpenStudy (anonymous):

:D